Attempt 6 Assume that the complete combustion of one mole of glucose to carbon dioxide and water liberates 2870 kJ/mol (AG
Answer #1

Total energy per mole glucose = -2870 kJ/mol

The energy is converted with an efficiency of 45%

So useful energy is 45% of this total energy. That is

45% of 2870 = 45× 2870/100 = 1291.5 kJ

57 kJ energy is used by number of cycle= 1

1291.5 kJ energy is used by no. Of cycle= 1×1291.5/57

= 22.7 ≈ 22 or 23 cycle